# IPCGRL: Language-Instructed Reinforcement Learning for Procedural Level Generation

### v2

@@ Abstract
Recent research has highlighted the significance of natural language in enhancing the controllability of generative models. While various efforts have been made to leverage natural language for content generation, research on deep reinforcement learning (DRL) agents utilizing text-based instructions for procedural content generation remains limited. In this paper, we propose IPCGRL, an instruction-based procedural content generation method via reinforcement learning, which incorporates a sentence embedding model. IPCGRL fine-tunes task-specific embedding representations to effectively compress game-level conditions. We evaluate IPCGRL in a two-dimensional level generation task and compare its performance with a general-purpose embedding method. The results indicate that IPCGRL achieves up to a 21.4% improvement in controllability and a 17.2% improvement in generalizability for unseen instructions. Furthermore, the proposed method extends the modality of conditional input, enabling a more flexible and expressive interaction framework for procedural content generation.
@@ I Introduction
Procedural content generation (PCG) has become an essential technique in game development and artificial intelligence, enabling the automatic creation of diverse, scalable, and adaptive game environments with minimal manual effort. Several machine learning-based approaches have been developed using search-based methods, generative adversarial networks, and text generative models [1]. With recent advancements in deep reinforcement learning (DRL), it has become possible to train generative models in a data-free manner [2]. The DRL approach optimizes an agent’s policy using a reward function and introduces variations in the policy to achieve different objectives. This approach has shown promising results in content generation tasks.1112 Equal contribution 3 Corresponding author
Research on procedural content generation via reinforcement learning (PCGRL) has explored various topics, including numerically controllable content generation [3], experience-driven generation [4], applications in three-dimensional game environments [5], scalability issues [6], graph generation [7], multiplayer content [8] and reward function design methodologies [9, 10]. Typically, PCGRL methods define a specific goal as a reward function within reinforcement learning, allowing an RL agent to learn and generate content accordingly. However, since these methods rely on predefined numerical features, they struggle to adapt to complex or evolving objectives and have limited capacity to incorporate high-level human intent and intervention. Therefore, a more intuitive and flexible approach to content generation is needed that can overcome the constraints of traditional PCGRL.
@@
Natural language is a powerful tool for controlling DRL agents, as it conveys human knowledge intuitively and serves as an effective interface. Language-instructed DRL enhances policy learning by leveraging natural language instructions, enabling agents to acquire goal-directed behaviors in domains such as robotics [11, 12] and multi-agent cooperation [13, 14]. However, despite these advantages, research on utilizing natural language as a conditioning input in RL-based content generation remains largely unexplored. Existing PCGRL approaches rely on predefined reward functions and manually designed constraints, which limit adaptability and expressiveness. Leveraging language-instructed approaches, the model can be more flexibly controlled through a user-friendly modality.
@@ II-A Controllable PCGRL
Controllable PCGRL is a technique to generate desired content with vectorized conditions. The authors of [3] first introduced the conditional DRL models in the two-dimensional level generation task. The conditional observation (e.g., path length) is defined as $c=\text{sign}(g-f(s_{t}))$, where $g$ is the goal condition and $f(s_{t})$ is the current metric of each condition. The conditional observation is concatenated to the observation of the one-hot level array. The loss of metric is calculated to $l_{t}=||g-s_{t}||_{L_{1}}$ and the agent gets a reward $r_{t}=l_{t-1}-l_{t}$. The previous studies [3, 5, 8] employ numerical conditional observation that there are restrictions on input modality.
@@ II-B Language-Instructed DRL
Language-instructed DRL is a methodology in which an agent learns to execute goal-directed behavior based on given instructions. Unlike traditional reinforcement learning, where the agent receives only the environmental state as input, this approach provides both the environmental state and a natural language instruction. This approach is widely utilized in various domains such as robotics [11, 12] and multi-agent cooperation [13, 14], and enables effective exploration and policy generalization in real-world environments where rewards are sparse or complex through the guidance of natural language instructions. Furthermore, language-instructed DRL aims to enhance the efficiency and adaptability of interactions by allowing the agent to directly understand and execute human linguistic directives.
@@ III-A Environment
This study employs a popular two-dimensional level generation gym environment [2, 3] and we select the GPU-accelerated version [6]. The modified Dungeon problem features three game tiles, Empty [Uncaptioned image], Wall [Uncaptioned image], and Bat [Uncaptioned image], and we use the Turtle representation for the environment configuration. The observation comprises the positions of each tile type, along with the location of the tile to be modified. The discrete action space includes seven behaviors: four actions correspond to moving the agent to the modification location, while the remaining three determine the specific tile type that replaces the tile at the modification location. The level size was set to $16\times 16$, and the episode length is configured to 1,500 steps per episode. Each episode begins with a randomly initialized matrix derived from the three-tile set.
@@ IV The IPCGRL Framework
We propose IPCGRL, a language-instructed PCGRL framework designed to train a language-instructed DRL agent for procedural level generation. Fig. 6 illustrates the two-phase training process to develop a language-instructed PCGRL agent. In the first phase, the encoder model is trained, and the trained encoder is then frozen during the training of the DRL agent in the second phase. The following sections describe the procedure for collecting instruction data (Section IV-A), training a language encoder model to compress conditional embeddings (Section IV-B), and training a DRL agent using compressed embedding features (Section IV-C). The source code and the instruction sets are available in this repository [^2] https://github.com/bic4907/insturcted_pcgrl.
@@ IV-B Conditional Instruction Encoder
The objective of the training instruction encoder proposed in this study is to generate an embedding vector for a given natural language instruction $\mathcal{I}$, denoted as $z_{enc}=[w_{1},w_{2},...,w_{n}]_{1\times{}n}$, where each vector component $w_{i}\in{}\mathbb{R}^{n}$ and $n\in{}\mathbb{N}$ represents the embedding dimension. Each natural language instruction $\mathcal{I}$ expresses a set of one or more task-condition pairs $(\tau_{i},c_{i})$. We define a function $\mathcal{C}$ that maps $\mathcal{I}$ to the set of task-condition pairs, where $\mathcal{C}(\mathcal{I})=\{(\tau_{\scriptscriptstyle\text{RG}},c_{\scriptscriptstyle\text{RG}}),...,(\tau_{\scriptscriptstyle\text{BD}},c_{\scriptscriptstyle\text{BD}})\}$. The encoder generates the embedding vector $z_{enc}=E_{\theta}(z_{bert})$, where the input vector is obtained as $z_{bert}=BERT_{\phi}(I)$, where $\phi$ is fixed. The decoder is trained to predict the fitness value as $f_{pred}=D_{\theta}(s^{\prime},z_{enc}),s^{\prime}\sim\mathcal{B}$, where state $s^{\prime}$ is sampled from the offline state dataset $\mathcal{B}$. The label fitness value is computed using the fitness function as follows Eq. 1.
@@ V-A Model Training Setting
We employ the pre-trained BERT model [^3] https://huggingface.co/docs/transformers/model_doc/bert[24] as the sentence embedding model, and stack two dense layers for both the encoder and decoder networks. The latent vector size of $z_{enc}$ is set to 64, i.e., $\mathbb{R}^{64}$. We collect a state buffer of 86K samples using the controllable PCGRL agent [6] to cover a diverse distribution of states, excluding duplicate state data from the dataset. The encoder and decoder models are trained for 100 epochs, and the fitness of the tasks is normalized to the range [-5, 5].
The DRL models are trained using proximal policy optimization (PPO) [25] for a total of 50 million timesteps, using the PureJaxRL implementation [26]. The hyperparameters are set as follows: $10$ epoch size, $128$ rollout length, $\gamma_{\text{GAE}}=0.95$, $\gamma=0.99$, $1^{-4}$ learning rate. Each comparison is evaluated over three independent runs, and the results are averaged. All experiments are conducted on machines equipped with a RTX 8000 GPU. The reward function weights are set to $w_{PL}=1$ and $w_{BC}=0.083$, considering reward magnitude and sparsity.
@@ V-B Evaluation Criteria
Progress ($P$) measures the relative percentage change from the initial state toward the specified target conditions. We improved the progress function to account for cases where the agent’s generated outcome exceeded the goal, leading to increased loss: $1-\lvert{\frac{g-s_{T}}{g-s_{0}}}\lvert$ where $T$ is the terminal state and normalized the loss with initial state. This function measures the progress of modifications toward achieving the goal state while ensuring that values remain within the range $[0,100]$. Additionally, it considers scenarios where exceeding the target results in inefficiencies or unintended negative consequences.
@@ VI-A Single-Task Generation with Encoded Latent Space
We assess the controllability of different methods by using the scalar-controlled generator as a reference point and primarily focus on comparing text-modality approaches. Fig. 7 illustrates the performance progress of five single-task instructions (Section IV-A), highlighting the differences between text- and scalar-conditioned generators. The generated levels are illustrated in Fig. 25.
@@ VI-B Generalizability on Unseen Instruction Set
The robustness to unseen instruction sentences is crucial for ensuring the generalizability of text-conditioned generative models. We evaluate model performance across three different instruction sets in the instruction dataset: training instructions within the same task (in-distribution, in-dist), test instructions within the same task but unseen during training (near out-of-distribution, near-OOD), and instructions from entirely different tasks (far-OOD). We additionally collect a near-OOD test set, comprising 25% of the number of training instructions, by modifying condition values through the replacement of quantifier words (e.g., many) that were not present in the training instruction set. We conduct the experiment on the PL task since its path structure is similar to the wall task (WC), but it uses different tile names, such as bat (BC, BD), and it provides the most neutral instruction among them.
Table II presents a comparative analysis of performance progression on seen and unseen instructions for $\tau_{\scriptscriptstyle\text{PL}}$. The scalar-controlled models were excluded from the near-OOD evaluation as the instruction split was determined by uniqueness. The IPCGRL ($z_{enc}$) model demonstrates superior performance compared to both the scalar-based and IPCGRL ($z_{bert}$) models under in-dist and near-OOD conditions, even outperforming scalar-controlled models in certain cases. The two IPCGRL methods underperform compared to other methods on far-OOD tasks, showing little generalization to other words included in different tasks. The task-specific latent representation ($z_{enc}$) achieves a 49.1% improvement on seen instructions and a 17.2% improvement on unseen instructions relative to the general-purpose embedding ($z_{bert}$). These findings indicate that task-specific embeddings enhance the model’s ability to generalize to unseen instructions within the same task domain. Moreover, this suggests that IPCGRL is capable of handling diverse numerical expressions during inference.

### v4

@@ Abstract
Recent research has highlighted the significance of natural language in enhancing the controllability of generative models. While various efforts have been made to leverage natural language for content generation, research on deep reinforcement learning (DRL) agents utilizing text-based instructions for procedural content generation remains limited. In this paper, we propose IPCGRL, an instruction-based procedural content generation method via reinforcement learning, which incorporates a sentence embedding model. IPCGRL fine-tunes task-specific embedding representations to effectively compress game-level conditions. We evaluate IPCGRL in a two-dimensional level generation task and compare its performance with a general-purpose embedding method. The results indicate that IPCGRL achieves up to a 21.4% improvement in controllability and a 17.2% improvement in generalizability for unseen instructions with varied condition expressions within the same task. Furthermore, the proposed method extends the modality of conditional input, enabling a more flexible and expressive interaction framework for procedural content generation.
@@ I Introduction
Procedural content generation (PCG) has become an essential technique in game development and artificial intelligence, enabling the automatic creation of diverse, scalable, and adaptive game environments with minimal manual effort. Several machine learning-based approaches have been developed using search-based methods, generative adversarial networks, and text generative models. With recent advancements in deep reinforcement learning (DRL), it has become possible to train generative models in a data-free manner [2]. The DRL approach optimizes an agent’s policy using a reward function and introduces variations in the policy to achieve different objectives. This approach has shown promising results in content generation tasks.
Research on procedural content generation via reinforcement learning (PCGRL) has explored various topics, including numerically controllable content generation [3], experience-driven generation [4], applications in three-dimensional game environments [5], scalability issues [6], multiplayer content [8] and reward function design methodologies [9, 10]. Typically, PCGRL methods define a specific goal as a reward function within reinforcement learning, allowing an RL agent to learn and generate content accordingly. However, since these methods rely on predefined numerical features, they struggle to adapt to complex or evolving objectives and have limited capacity to incorporate high-level human intent and intervention. Therefore, a more intuitive and flexible approach to content generation is needed that can overcome the constraints of traditional PCGRL.
@@
Natural language is a powerful tool for controlling DRL agents, as it conveys human knowledge intuitively and serves as an effective interface. Language-instructed DRL enhances policy learning by leveraging natural language instructions, enabling agents to acquire goal-directed behaviors in domains such as robotics [12] and multi-agent cooperation [13]. However, despite these advantages, research on utilizing natural language as a conditioning input in RL-based content generation remains largely unexplored. Existing PCGRL approaches rely on predefined reward functions and manually designed constraints, which limit adaptability and expressiveness. Leveraging language-instructed approaches, the model can be more flexibly controlled through a user-friendly modality.
@@ II-A Controllable PCGRL
Controllable PCGRL is a technique to generate desired content with vectorized conditions. The authors of [3] first introduced the conditional DRL models in the two-dimensional level generation task. The conditional observation (e.g., path length) is defined as $c=\text{sign}(g-f(s_{t}))$, where $g$ is the goal condition and $f(s_{t})$ is the current metric of each condition. The conditional observation is concatenated to the observation of the one-hot level array. The loss of metric is calculated to $l_{t}=||g-f(s_{t})||_{L_{1}}$ and the agent gets a reward $r_{t}=l_{t-1}-l_{t}$. The previous studies [3, 5, 8] employ numerical conditional observation that there are restrictions on input modality.
@@ II-B Language-Instructed DRL
Language-instructed DRL is a methodology in which an agent learns to execute goal-directed behavior based on given instructions. Unlike traditional reinforcement learning, where the agent receives only the environmental state as input, this approach provides both the environmental state and a natural language instruction. This approach is widely utilized in various domains such as robotics [12] and multi-agent cooperation [13]. It enables effective exploration and policy generalization in real-world environments–where rewards are sparse or complex–through the guidance of natural language instructions. Furthermore, language-instructed DRL aims to enhance the efficiency and adaptability of interactions by allowing the agent to directly understand and execute human linguistic directives.
@@ III-A Environment
This study employs a popular two-dimensional level generation gym environment [2, 3] and we select the GPU-accelerated version [6]. The modified Dungeon problem features three game tiles: Empty [Uncaptioned image], Wall [Uncaptioned image], and Bat [Uncaptioned image]; we use the Turtle representation, which the action space incorporates relative agent movement and tile replacement. The observation comprises the positions of each tile type, along with the location of the tile to be modified. The discrete action space includes seven behaviors: four actions correspond to moving the agent to the modification location, while the remaining three determine the specific tile type that replaces the tile at the modification location. The level size was set to $16\times 16$, and the episode length is configured to 1,500 steps per episode. Each episode begins with a randomly initialized matrix derived from the three-tile set.
@@ IV The IPCGRL Framework
We propose IPCGRL, a language-instructed PCGRL framework designed to train a language-instructed DRL agent for procedural level generation. Fig. 6 illustrates the two-phase training process to develop a language-instructed PCGRL agent. In the first phase, the encoder model is trained, and the trained encoder is then frozen during the training of the DRL agent in the second phase. The following sections describe the procedure for collecting instruction data (Section IV-A), training a language encoder model to compress conditional embeddings (Section IV-B), and training a DRL agent using compressed embedding features (Section IV-C). The source code and the instruction sets are available in this repository [^1] https://github.com/bic4907/language-instructed-pcgrl.
@@ IV-B Conditional Instruction Encoder
The objective of the training instruction encoder proposed in this study is to generate an embedding vector for a given natural language instruction $\mathcal{I}$. This embedding, denoted as $z_{enc}$, is a 64-dimensional vector that captures the semantic representation of the instruction. The encoder generates the embedding vector $z_{enc}=E_{\theta}(z_{bert})$, where the input vector is obtained as $z_{bert}=BERT_{\phi}(\mathcal{I})$. Here, $\phi$ denotes the fixed pretrained parameters of the BERT model. The decoder is trained to predict the fitness value as $f_{pred}=D_{\theta}(s^{\prime},z_{enc}),s^{\prime}\sim\mathcal{B}$, where state $s^{\prime}$ is sampled from the offline state dataset $\mathcal{B}$. The label fitness value is computed using the fitness function as follows Eq. 1.
@@ V-A Model Training Setting
We employ the pre-trained BERT model [24] as the sentence embedding model, and stack two dense layers for both the encoder and decoder networks. The BERT output vector $z_{\text{bert}}\in\mathbb{R}^{768}$ is projected into a latent space $z_{\text{enc}}\in\mathbb{R}^{64}$ through the encoder network. We collect a state buffer of 86K samples using the controllable PCGRL agent [6] to cover a diverse distribution of states, excluding duplicate state data from the dataset. Since the RL policy takes both the state and the condition as input, it is crucial to learn an embedding representation that generalizes well to the state distribution. The encoder and decoder models are trained for 100 epochs, and the fitness of the tasks is normalized to the range [-5, 5].
The DRL models are trained using proximal policy optimization (PPO) [25] for a total of 50 million timesteps, using the PureJaxRL implementation [26]. The chosen timesteps setting ensures sufficient convergence time in both single- and multi-task experiments. The hyperparameters are set as follows: $10$ epoch size, $128$ rollout length, $\gamma_{\text{GAE}}=0.95$, $\gamma=0.99$, $10^{-4}$ learning rate. Each comparison is evaluated over three independent runs, and the results are averaged. All experiments are conducted on machines equipped with a RTX 8000 GPU. The weights of the reward functions are set to $w_{PL}=1$ and $w_{BC}=0.083$, considering reward magnitude and sparsity for the multi-task experiment.
@@ V-B Evaluation Criteria
Progress ($P$) measures the relative percentage change from the initial state toward the specified target conditions. We improved the progress function to account for cases where the agent’s generated outcome exceeded the goal, leading to increased loss: $1-\lvert{\frac{g-s_{T}}{g-s_{0}}}\lvert$ where $s_{T}$ is the terminal state, and the loss is normalized with respect to the initial state $s_{0}$. This function measures the progress of modifications toward achieving the goal state using a normalized score, which is scaled by 100 to produce a value within the range $[0,100]$. Additionally, it considers scenarios where exceeding the target results in inefficiencies or unintended negative consequences.
@@ VI-A Single-Task Generation with Encoded Latent Space
We assess the controllability of different methods by using the scalar-controlled generator as a reference point and primarily focus on comparing text-modality approaches. Fig. 7 illustrates the performance progress of five single-task instructions (Section IV-A), highlighting the differences between text- and scalar-conditioned generators. The generated levels are illustrated in Fig. 25. To compare the embedding representation capability within the same modality, we compute an improvement metric by normalizing the performance of the proposed IPCGRL ($z_{\text{enc}}$) with respect to the baseline IPCGRL ($z_{\text{bert}}$).
@@ VI-B Generalizability on Unseen Instruction Set
The robustness to unseen instruction sentences is crucial for ensuring the generalizability of text-conditioned generative models. We evaluate model performance across three different instruction sets in the instruction dataset: training instructions within the same task (in-distribution, in-dist), test instructions within the same task but unseen during training (near out-of-distribution, near-OOD), and instructions from entirely different tasks (far-OOD). We construct a near-OOD test set comprising 25% of the number of training instructions by modifying condition values through the replacement of quantifier words (e.g., many). The average cosine similarity measured with BERT embeddings is $0.831\pm 0.084$ between the train- and test instructions, indicating that the test set retains task semantics while introducing meaningful lexical variation. We conduct the experiment on the PL task since its path structure is similar to the wall task (WC), but it uses different tile names, such as bat (BC, BD), and it provides the most neutral instruction among them.
Table II presents a comparative analysis of performance progression on seen and unseen instructions for $\tau_{\scriptscriptstyle\text{PL}}$. The scalar-controlled models were excluded from the near-OOD evaluation as the instruction split was determined by uniqueness. Although the three evaluation sets are similar, slight differences in their conditions necessitate comparing methods only within the same set to ensure fairness. The IPCGRL ($z_{enc}$) model demonstrates superior performance compared to both the scalar-based and IPCGRL ($z_{bert}$) models under in-dist and near-OOD conditions, even outperforming scalar-controlled models in certain cases. The two IPCGRL methods underperform compared to other methods on far-OOD tasks, showing little generalization to other words included in different tasks. The task-specific latent representation ($z_{enc}$) achieves a 49.1% improvement on seen instructions and a 17.2% improvement on unseen instructions relative to the general-purpose embedding ($z_{bert}$). These findings indicate that task-specific embeddings enhance the model’s ability to generalize to unseen instructions within the same task domain. Moreover, this suggests that IPCGRL is capable of handling diverse numerical expressions during inference.
@@ VIII Conclusion and Future Work
IX Acknowledgments
This research was supported by Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2025 (Project Number: RS-2024-00441523).
This research was also supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (RS-2024-00395665).